Correctly answer each part of question 2
2 Liquidity ratios Aa Aa Most firms borrow money to finance some of their assets, and most will choose to borrow some long-term funds and some short-term funds. Which group of lenders would put greater emphasis on a firms liquidity ratio when evaluating a potential borrower? Short-term lenders O Long-term lenders
follows: The most recent data from the annual balance sheets of N8B Equipment Company and Scramouche Opera Company are as Balance Sheet December 31st (Millions of dollars) Scramouche N&B Scramouche N&B Opera Equipment Company Company Opera Equipment Company Company Liabilities Current liabilities Assets Current assets $0 63 359 422 515 937 $184 Accounts payable $0 Cash Accounts receivable Inventories $287 105 308 700 68 Accruals 198 Notes payable 450 Total current liabilities Total current assets Net fixed assets Net plant and equipment 337 337 413 750 Long-term bonds 550 550 Total debt Common equity Common stock 203 110 313 1,250 163 87 250 1,000 Retained earnings Total common equity Total liabilities and equity Total assets 1,250 1,000
, and its quick ratio is Scramouche Opera Companys current ratio N&B Equipment Companys current ratio is is , and its quick ratio is . Note: Round your values to four decimal places. Which of the following statements are true? Check all that apply. N&B Equipment Company has less liquidity but also a greater reliance on outside cash flow to finance its short-term obligations than Scramouche Opera Company. A current ratio of 1 indicates that the book value of the companys current assets is equal to the book value of its current liabilities. An increase in the quick ratio over time usually means that the companys liquidity position is improving and that the company is managing its short-term assets well. N&B Equipment Company has a better ability to meet its short-term liabilities than Scramouche Opera Company An increase in the current ratio over time always means that the companys liquidity position is improving.
Answer #1

Short term lenders since their dues are dependent on liquidity of the company

Quick ratio = Quick Assets/Current Liabilities

Current ratio = Current Assets/ Current Liabilities

Quick Assets = Current Assets - Inventory

Scramoucha Opera

Quick ratio = 392/422 = 0.9289

Current ratio = 700/422 = 1.6588

N&B

Quick Ratio = 252/337 = 0.7478

Current Ratio = 450/337 = 1.3353

First 3 statements are correct

Improvement of current ratio might also mean piling up of inventory which is not good